Sylvia wants to lose 40 pounds. How many kcalories does this represent?

Group of answer choices
105,000
85,000
140,000
3500
Flag this Question
Question 2 1 pts
0multiple_choice_question 97263028
If Paul eats 350 kcals less every day, how many kcalories does he not consume each week?
If Paul eats 350 kcals less every day, how many kcalories does he not consume each week?
Group of answer choices
2450
4000
3500
2280
Flag this Question
Question 3 1 pts
0multiple_choice_question 97263111
How much weight could Paul expect to lose in two weeks if he has a calorie deficit of 700 kcals each day?
How much weight could Paul expect to lose in two weeks if he has a calorie deficit of 700 kcals each day?
Group of answer choices
3.5 pounds
1.4 pounds
2.8 pounds
4.2 pounds
Flag this Question
Question 4 1 pts
0multiple_choice_question 97263190
What is considered to be a safe rate of weight loss according to information presented in class?
What is considered to be a safe rate of weight loss according to information presented in class?
Group of answer choices
none of the answers given
20% loss of original body weight in one month
up to 2 pounds per week
less than 5 pounds per week
Flag this Question
Question 5 2 pts
0multiple_choice_question 97262808
Is Paul practicing a safe rate of weight loss with his 600 kcal deficit each day? Why or why not?
Is Paul practicing a safe rate of weight loss with his 600 kcal deficit each day? Why or why not?
Group of answer choices
Yes, he is only losing 1.2 pounds per week.
Yes, he is losing just a little over two pounds per week.
No, he is losing 2.4 pounds per week.
No, he is losing too much.
Answer #1

1)

3500 kcal = 1 pound

So 40 pounds = 40*3500 = 140000 kcal

2)

Every day he doesn't eat 350 kcal.

In 7 days = 350 * 7 = 2450 kcal

3)

Per day = 700 kcal

In 2 weeks = 7000 * 14 = 9800 kcal

3500 kcal = 1 pound

9800 kcal = 9800/3500 = 2.8 pounds

4)

The safe rate of weight loss is 1-2 pounds per week.

5)

600 kcal per day

Per week 600*7 =4200 kcal

3500 kcal =1 pound

4200 kcal = 4200/3500 = 1.2 pounds per week

So answer is A.